# Notes on the Practical Application of Nested Sampling: MultiNest, (non)convergence, and rectification

### v1

@@ 4.2 Abnormal Distributions I: Bananas           (Generalized Rosenbrock Distributions)
  • MultiNest struggles in higher dimensions more so than when sampling simpler normal distributions of the same dimensionality, requiring $\epsilon\leq 0.01$ at $N_{\rm live}\sim 100$ and $\epsilon\leq 0.1$ for $N_{\rm live}\sim 10^{4}$ when sampling a 10D distribution.
The generalized Rosenbrock distribution evidence estimates are presented in Figure 4. Overall, I have found that MultiNest is able to sample from generalized Rosenbrock distributions without too much additional difficulty compared to an uncorrelated multivariate normal distributions. For a given dimensionality, smaller sampling efficiency values are required to achieve converged results, in comparison to sampling simpler normal distributions.
@@ 4.3 Abnormal Distributions II: Log-Gammas
  • MultiNest is slightly biased towards underestimating the widths of posteriors.
The above observations are largely drawn from Figures 6 and 7. The former figure presents the individual evidence estimates from each multivariate normal analysis; although error bars are technically included, the estimated errors are less than unity and completely dwarfed by the systematic errors present in most of the analyses. In most of the cases where MultiNest does a very poor job of estimating the evidence, it also derives very incorrect posterior estimates, reporting results with errors on the order of $100\%$. If this holds in general, consistent evidence measurements might suggest fairly trustworthy posterior inferences.
A more worrying aspect of Figures 6 and 7 is the breadth over which MultiNest can report $\sim 100\%$ erroneous results. Let us take the 10-dimensional problem as an example. Based on the recommendations in the MultiNest documentation, one might reasonably start with 1000 live points and an efficiency of 0.3, which would almost certainly lead to a log-evidence error on the order of $\sim 500.$ One may need to decrease the sampling efficiency by a factor of $\sim 300$ holding the number of live points fixed, or by a factor of $\sim 3-100$ while increasing the number of live points by factors of 10-3 to derive accurate results. However, one can not know this a priori, so diligent convergence testing should always be practiced, until the evidence estimates and posteriors converge.
Again measuring the widths of the analytical and measured ±$\pm 1\sigma$ credible intervals and restricting the samples to only those with $L_{1}$ errors less than 0.2, 2110 out of 4008 estimated credible regions were too narrow. This is again quite improbable (if MultiNest underestimated the width of credible regions as often as it overestimated them, there would be a $\sim 0.04\%$ chance of this many or more of that sample being too narrow), although the evidence is not as strong given the smaller data set. However, looking at the full set of posterior estimates, I found that 3350 out of 7650 credible interval widths were underestimated. Here there is a silver lining: when MultiNest does especially poorly, credible interval widths tend to be overestimated, at least for this problem.
@@ 5 MultiNest in Action:
More specifically, at larger values of the sampling efficiency (such as $\epsilon=0.3,$ which is still less than half that recommended by the MultiNest documentation for parameter estimation), MultiNest produces systematically narrow posterior estimates. These biases do not disappear as more live points are used in the analyses; rather, as live points are added the run-to-run variation seems to decrease while the median error remains roughly constant between runs. The biases are significantly reduced at $\epsilon=0.1,$ and virtually disappear for this problem by $\epsilon\leq 0.03$. Intriguingly, for this particular problem, results can become slightly more biased as tol, the evidence tolerance used by MultiNest as a stopping condition, is decreased. Presumably, because the actual likelihood surface is nonsmooth, MultiNest’s bounding ellipsoids have a tendency to systematically underestimate the actual isolikelihood surface at a given iteration. Since the noise introduced in this test is a perturbation overtop a general trend, once MultiNest’s bounding hyperellipsoids are enlarged sufficiently at each stage they are able to derive much less biased samples. This tendency should hold for fairly generic, even smooth, likelihoods, but the this particular likelihood function seems to exacerbate the issue.
@@ 6 Strategies for Improvement
For problems like those analyzed in the present study, which require only a modest personal computer, a sound strategy, besides conducting rigorous convergence tests, is to use a more robust nested sampling tool. A number of other tools seem to meet the criterion of being less prone to biases than MultiNest, including POLYCHORD (Handley et al., 2015; Lemos et al., 2023), JAXNS (Albert, 2020), dynesty (Speagle, 2020), or UltraNest (Buchner, 2021). [^8] Empirical evidence points to slice sampling implementations being less prone to producing biased evidence estimates than the rejection sampling-based MultiNest (see, for example, Lemos et al. (2023) and Figure 4 of Albert (2020), although in that case other parameters must be tested to ensure convergence); other packages such as UltraNest implement more robust bounding regions for rejection sampling (e.g., Buchner, 2016a, 2021). Furthermore, packages such as dynesty and dyPolyChord implement dynamic nested sampling (Higson et al., 2019b), which can target both posterior and evidence accuracy in addition to providing the capability of further refining an initial nested sampling analysis; UltraNest also allows further refinements following an initial run, and targeting both posterior and evidence accuracy. If a given likelihood function is compatible with the JAX framework (Bradbury et al., 2018), JAXNS might be particularly favorable, and significantly faster than the alternatives (Albert, 2020). [^9] UltraNest also provides vectorized step samplers that can take advantage of JAX-based or otherwise GPU-parallelized likelihoods. Otherwise, MultiNest can remain an order of magnitude or so faster than more robust sampling algorithms. As shown above, one might need to improve the resolution (and possibly expense) of a fiducial MultiNest run by orders of magnitude for the results to be worth trusting, but both this prospect and that of using a slower sampling strategy can be difficult pills to swallow when committing thousands or millions of core hours to a Bayesian analysis on a supercomputer.
@@
Generally speaking, it is possible to use an initial (presumably relatively quick) MultiNest analysis to accelerate other methods more capable of accurately estimating evidences and posteriors. One approach is to use the initial nested sampling analysis to initialize a Markov-chain Monte Carlo (MCMC) analysis (e.g., Miller et al., 2019, 2021; Wolfe et al., 2023), where as long as the initial nested sampling analysis was able to correctly identify the most significant mode within a given parameter space the MCMC analysis can be expected to draw unbiased samples from the posteriors from the target distribution. If the initial nested sampling analysis was able to adequately sample the posterior distribution, the MCMC sampler will be able to start drawing unbiased samples almost instantly. However, if the nested sampling algorithm does a poor job of estimating posteriors this method can require a long “burn-in” period, and standard MCMC techniques can not reliably estimate the evidence. [^10] Parallel-tempered MCMC samples may be used to estimate the evidence as well (e.g., Gelman & Meng, 1998; Vousden et al., 2016), but estimates using only samples of the posterior typically have infinite variance (Newton & Raftery, 1994) and are therefore useless. See Wolfe et al. (2023) for a description of how samples from a single nested sampling analysis can be used to initialize a parallel-tempered MCMC analysis. Another related approach, although I have not yet tested this personally, could be to initialize a population Monte Carlo analysis (e.g., Cappé et al., 2004, 2008), similar to the description in Beaujean & Caldwell (2013), which could also produce evidence estimates via importance sampling.
Another strategy is to use a preliminary MultiNest analysis to jump-start a nested sampling analysis with a slower but more robust nested sampling strategy. Qualitatively, information about the posterior can be used to reduce the time spent by a nested sampling analysis compressing from the prior to the posterior (Petrosyan & Handley, 2022). This strategy can be used to refine a dubious initial analysis, or to accelerate the analysis a problem very similar to one already analyzed (for example, one adding a few new observations to an existing data set). Here, I’ve used the implementation of “warm starting” built into UltraNest. [^11] https://johannesbuchner.github.io/UltraNest/example-warmstart.html. One must take care when starting from MultiNest outputs, which have a slightly different format that UltraNest outputs.
@@ 6.1 MCMC
For this test, I began with arguably the most difficult problem — the 10D log-gamma distribution — and started with a MultiNest analysis that missed the mark, one of the $N_{\rm live}=1000,\,\epsilon=0.1$ runs. I first performed a Gaussian kernel density estimate of the MultiNest posterior samples; at this stage I prefer to use a bandwidth an order of magnitude or two (in this case a factor of 10) smaller than the Silverman suggestion (Silverman, 1986) to help prevent the kernel density estimate from extending beyond the prior bounds, if applicable. After estimating the density of the posterior, I resampled it, drawing 1024 samples, which I used to initialize the walkers in an MCMC analysis. For this I employed emcee version 3.1.4, using the default settings (Foreman-Mackey et al., 2013).
Some of the results from this test (for every other parameter) are shown in Figure 9. In this Figure, I’ve plotted the probability distributions derived from the first, second, and third groups of 750 MCMC iterations, after initializing the walkers based on draws from the MultiNest-derived posteriors. Over time, the walkers move from their initial positions to instead match the posterior. This sort of behavior can be expected as long as MultiNest is able to identify the most prominent mode in a multimodal parameter space. Because selecting good initial walker positions can be one of the challenges in an MCMC analysis, these two methods can complement each other nicely.
@@ 6.2 Starting Warmly
For this test, I began with the same MultiNest outputs as Section 6.1, but I instead used it to warm-start an UltraNest analysis. This technique has already been demonstrated for cases where the information used to warm-start a given nested sampling run is actually representative of the underlying posterior (e.g. Petrosyan & Handley, 2022). In this sense, the present test, where the nested sampling run used to warm-start the analysis was quite far from representing the posterior, is something like a worst-case scenario, and demonstrates the robustness of this strategy when the deviations between the warm start and actual posterior are large.
Despite using far from optimal samples to inform the warm-starting procedure, UltraNest was still able to perform adequately without any tuning. For example, UltraNest derived an evidence estimate with an error of $\Delta\log{Z}=0.690\pm 0.605$. On the other hand, MultiNest had an evidence error of $\Delta\log{Z}=-297.245\pm 0.259$. Additionally, as illustrated in Figure 10, UltraNest was able to derive much more accurate posterior estimates.
@@ 7 Conclusions
  • can derive completely erroneous posterior estimates,
  • is more likely to underestimate the width of a credible interval than to overestimate it.
The degree to which MultiNest produces biased results (or whether it instead produces good results) depends sensitively on the dimensionality and structure of a given problem.
Additionally, when sampling noisy likelihoods, MultiNest tends to produce posterior estimates that are systematically narrow; this bias can be ameliorated by decreasing the sampling efficiency, but is largely independent of the number of live points used. Although the particular test problem examined here is not necessarily analogous to practical applications, this behavior mimics that observed in some actual analyses where MultiNest seemed to significantly underestimate the widths of credible regions (e.g., Miller et al., 2019; Ih & Kempton, 2021; Miller et al., 2021). Because MultiNest’s posterior estimates can be systematically biased at larger sampling efficiency, it is crucial that any analysis relying on MultiNest-derived posteriors test convergence with respect to the sampling efficiency.
@@ Acknowledgments
This work benefited from stimulating discussions with and useful feedback from Arjun Savel, Jegug Ih, Matt Nixon, and Cole Miller. I am grateful for the detailed feedback and comments provided by Johannes Buchner on an draft of this manuscript. This work was supported in part by NASA ADAP grants 80NSSC21K0649 and 80NSSC20K0288. Many of these calculations were performed on the ASTRA computing cluster, which is maintained by the Department of Astronomy at the University of Maryland. I also used NASA’s Astrophysics Data System Bibliographic Services and the ArXiv extensively.

### v2

@@ 4.2 Abnormal Distributions I: Bananas           (Generalized Rosenbrock Distributions)
  • In the 10-dimensional case, MultiNest struggles to produce accurate results more so than when sampling simpler normal distributions, requiring $\epsilon\leq 0.01$ at $N_{\rm live}\sim 100$ and $\epsilon\leq 0.1$ for $N_{\rm live}\sim 10^{4}$.
The generalized Rosenbrock distribution evidence estimates are presented in Figure 4. Overall, I have found that MultiNest is able to sample from generalized Rosenbrock distributions without too much additional difficulty compared to an uncorrelated multivariate normal distributions. Because the isolikelihood surfaces deviate from hyperellipsoinds, for a given dimensionality smaller sampling efficiency values are required to achieve converged results, in comparison to sampling simpler normal distributions.
@@ 4.3 Abnormal Distributions II: Log-Gammas
The above observations are largely drawn from Figures 6 and 7. The former figure presents the individual evidence estimates from each multivariate log-gamma analysis; although error bars are technically included, the estimated errors are less than unity and completely dwarfed by the systematic errors present in most of the analyses. In most of the cases where MultiNest does a very poor job of estimating the evidence, it also derives very incorrect posterior estimates, reporting results with errors on the order of $100\%$. If this holds in general, consistent evidence measurements might suggest fairly trustworthy posterior inferences. However, as shown in Section 4.1, accurate posteriors do not imply accurate evidence measurements.
A more worrying aspect of Figures 6 and 7 is the breadth over which MultiNest can report $\sim 100\%$ erroneous results. Let us take the 10-dimensional problem as an example. Based on the recommendations in the MultiNest documentation, one might seemingly conservatively start with 1000 live points and an efficiency of 0.3, which would lead to a log-evidence error on the order of $\sim 500.$ One may need to decrease the sampling efficiency by a factor of $\sim 300$ holding the number of live points fixed, or by a factor of $\sim 3-100$ while increasing the number of live points by factors of 10-3 to derive accurate results. However, one can not know this a priori, so diligent convergence testing should always be practiced, until the evidence estimates and posteriors converge.
As in Section 4.2, MultiNest was somewhat challenged by isolikelihood surfaces that strongly deviate from hyperellipsoids, as illustrated in Figure 5 for the two-dimensional case. Over a wide range of hyperparameter settings, MultiNest derived posteriors with order-unity errors, the cause of which is illustrated in Figure 5: when too few live points or too large a sampling efficiency are used, MultiNest often misses the peak of the likelihood distribution entirely. Because the tail of the log-gamma distribution is much more shallow than the Gaussian or Rosenbrock distributions, it is much easier for all of the live points to end up in the tail with none sampling the peak. Clearly, using more live points makes it less likely to miss the peak entirely, and a smaller sampling efficiency safeguards against such situations.
@@ 5 MultiNest in Action:
More specifically, at larger values of the sampling efficiency (such as $\epsilon=0.3,$ which is still less than half that recommended by the MultiNest documentation for parameter estimation), MultiNest produces systematically narrow posterior estimates. These biases do not disappear as more live points are used in the analyses; rather, as live points are added the run-to-run variation seems to decrease while the median error remains roughly constant between runs. The biases are significantly reduced at $\epsilon=0.1,$ and virtually disappear for this problem by $\epsilon\leq 0.03$. Intriguingly, for this particular problem, results can become slightly more biased as tol, the evidence tolerance used by MultiNest as a stopping condition, is decreased. Presumably, because the actual likelihood surface is nonsmooth, MultiNest’s bounding ellipsoids have a tendency to systematically underestimate the actual isolikelihood surface at a given iteration. Unlike the smooth test problems considered earlier, the likelihood surface in this test was highly non-monotonic and multimodal; in this case, increasing the number of live points enough to adequately sample every mode present at a given log likelihood is unfeasible, but decreasing the sampling efficiency is tractable. Since the noise introduced in this test is a perturbation overtop a general trend, once MultiNest’s bounding hyperellipsoids are enlarged sufficiently at each stage they are able to draw much less biased samples.
@@ 6 Strategies for Improvement
For problems like those analyzed in the present study, which require only a modest personal computer, a sound strategy, besides conducting rigorous convergence tests, is to use a more robust nested sampling tool. A number of other tools seem to meet the criterion of being less prone to biases than MultiNest, including POLYCHORD (Handley et al., 2015; Lemos et al., 2023), JAXNS (Albert, 2020), dynesty (Speagle, 2020), and UltraNest (Buchner, 2021). [^11] Empirical evidence points to slice sampling implementations being less prone to producing biased evidence estimates than the rejection sampling-based MultiNest (see, for example, Lemos et al. (2023) and Figure 4 of Albert (2020), although in that case other parameters must be tested to ensure convergence); other packages such as UltraNest implement more robust bounding regions for rejection sampling (e.g., Buchner, 2016a, 2021). Furthermore, packages such as dynesty and dyPolyChord implement dynamic nested sampling (Higson et al., 2019b), which can target both posterior and evidence accuracy in addition to providing the capability of further refining an initial nested sampling analysis; UltraNest also allows further refinements following an initial run, and targeting both posterior and evidence accuracy. If a given likelihood function is compatible with the JAX framework (Bradbury et al., 2018), JAXNS might be particularly favorable, and significantly faster than the alternatives (Albert, 2020). [^12] UltraNest also provides vectorized step samplers that can take advantage of JAX-based or otherwise GPU-parallelized likelihoods. Otherwise, MultiNest can remain an order of magnitude or so faster than more robust sampling algorithms. As shown above, one might need to improve the resolution (and possibly expense) of a fiducial MultiNest run by orders of magnitude for the results to be worth trusting, but both this prospect and that of using a slower sampling strategy can be intractable when committing thousands or millions of core hours to a Bayesian analysis on a supercomputer.
Generally speaking, it is possible to use an initial (presumably relatively quick) MultiNest analysis to accelerate other methods more capable of accurately estimating evidences and posteriors. One approach is to use the initial nested sampling analysis to initialize a Markov-chain Monte Carlo (MCMC) analysis (e.g., Miller et al., 2019, 2021; Wolfe et al., 2023), where as long as the initial nested sampling analysis was able to correctly identify the most significant mode within a given parameter space the MCMC analysis can be expected to draw unbiased samples from the posterior. If the initial nested sampling analysis was able to adequately sample the posterior distribution, the MCMC sampler will be able to start drawing unbiased samples almost instantly. However, if the nested sampling algorithm does a poor job of estimating posteriors this method can require a long “burn-in” period, and standard MCMC techniques can not reliably estimate the evidence. [^13] Parallel-tempered MCMC samples may be used to estimate the evidence as well (e.g., Gelman & Meng, 1998; Vousden et al., 2016), but estimates using only samples of the posterior typically have infinite variance (Newton & Raftery, 1994) and are therefore useless. See Wolfe et al. (2023) for a description of how samples from a single nested sampling analysis can be used to initialize a parallel-tempered MCMC analysis. Another related approach, although I have not yet tested it personally, could be to initialize a population Monte Carlo analysis (e.g., Cappé et al., 2004; Cappé2008), similar to the description in Beaujean & Caldwell (2013), which could also produce evidence estimates via importance sampling.
Another strategy is to use a preliminary MultiNest analysis to jump-start a nested sampling analysis with a slower but more robust nested sampling strategy. Qualitatively, information about the posterior can be used to reduce the time spent by a nested sampling analysis compressing from the prior to the posterior (Petrosyan & Handley, 2022). This strategy can be used to refine a dubious initial analysis, or to accelerate the analysis a problem very similar to one already analyzed (for example, one adding a few new observations to an existing data set). Here, I have used the implementation of “warm starting” built into UltraNest. [^14] https://johannesbuchner.github.io/UltraNest/example-warmstart.html. One must take care when starting from MultiNest outputs, which have a slightly different format that UltraNest outputs.
@@ 6.1 MCMC
For this test, I began with arguably the most difficult problem — the 10D log-gamma distribution — and started with a MultiNest analysis that had used inadequate hyperparameters, one of the $N_{\rm live}=1000,\,\epsilon=0.1$ runs. I first performed a Gaussian kernel density estimate of the MultiNest posterior samples; at this stage I prefer to use a bandwidth an order of magnitude or two (in this case a factor of 10) smaller than the Silverman suggestion (Silverman, 1986) to help prevent the kernel density estimate from extending beyond the prior bounds, if applicable. After estimating the density of the posterior, I resampled it, drawing 1024 samples, which I used to initialize the walkers in an MCMC analysis. For this I employed emcee version 3.1.4, using the default settings (Foreman-Mackey et al., 2013).
Some of the results from this test (for every other parameter) are shown in Figure 9. In this Figure, I have plotted the probability distributions derived from the first, second, and third groups of 750 MCMC iterations, after initializing the walkers based on draws from the MultiNest-derived posteriors. Over time, the walkers move from their initial positions to instead match the posterior. This sort of behavior can be expected as long as MultiNest is able to identify the most prominent mode in a multimodal parameter space. This procedure also necessitates diligence during the MCMC stage to ensure convergence; as an example, if this analysis had used too few walkers, a less-efficient MCMC proposal, or fewer iterations, accurate posteriors would not have been recovered. Because selecting good initial walker positions can be one of the greatest challenges in an MCMC analysis, these two methods can complement each other nicely.
@@ 6.2 Starting Warmly
For this test, I began with the same MultiNest outputs as Section 6.1, but I instead used it to warm-start an UltraNest analysis. This technique has already been demonstrated for cases where the information used to warm-start a given nested sampling run is representative of the underlying posterior (e.g. Petrosyan & Handley, 2022). In this sense, the present test, where the nested sampling run used to warm-start the analysis did not accurately measure the posterior, is something like a worst-case scenario, and demonstrates the robustness of this strategy when the deviations between the warm start and actual posterior are large.
Despite using suboptimal samples to inform the warm-starting procedure, UltraNest was still able to perform adequately without any tuning. For example, UltraNest derived an evidence estimate with an error of $\Delta\log{Z}=0.690\pm 0.605$. On the other hand, MultiNest had an evidence error of $\Delta\log{Z}=-297.245\pm 0.259$. Additionally, as illustrated in Figure 10, UltraNest was able to in this case derive more accurate posterior estimates. Although the aforementioned UltraNest results are encouraging, this procedure should not be taken for granted, and one still ought to perform convergence tests as part of practical analyses.
@@ 7 Conclusions
  • can derive completely erroneous posterior estimates
The degree to which MultiNest produces biased results (or whether it instead produces accurate results) depends sensitively on the dimensionality and structure of a given problem. Even with relatively few ($\sim 100$) live points, MultiNest can still derive accurate evidence estimates and posteriors given a sampling efficiency value low enough so that at each stage it draws samples from within the correct isolikelihood surface. MultiNest naturally experiences more difficulty sampling from likelihoods that are more difficult to approximate using ellipsoids (such as in Sections 4.2 and 4.3); since this is rarely known a priori, testing convergence is a necessary part of any analysis.
Additionally, when sampling noisy likelihoods, MultiNest tends to produce posterior estimates that are systematically narrow; this bias can be ameliorated by decreasing the sampling efficiency, but is largely independent of the number of live points used. Although the particular test problem examined here is not necessarily analogous to any practical application, this behavior mimics that observed in some actual analyses where MultiNest seemed to significantly underestimate the widths of credible regions (e.g., Miller et al., 2019; Ih & Kempton, 2021; Miller et al., 2021). Because MultiNest’s posterior estimates can be systematically biased at larger sampling efficiencies, it is crucial that any analysis relying on MultiNest-derived posteriors test convergence with respect to the sampling efficiency, and not only the number of live points.
@@ Acknowledgments
This work benefited from stimulating discussions with and useful feedback from Arjun Savel, Jegug Ih, Matt Nixon, and Cole Miller. I am grateful for the detailed feedback and comments provided by Johannes Buchner on an draft of this manuscript. This work was supported in part by NASA ADAP grants 80NSSC21K0649 and 80NSSC20K0288. I thank the reviewers for their numerous comments that led to substantial improvements of this paper. Many of these calculations were performed on the ASTRA computing cluster, which is maintained by the Department of Astronomy at the University of Maryland. I also used arXiv and NASA’s Astrophysics Data System Bibliographic Services extensively.